A proton accelerates from rest in a uniform electric field of 615 N/C. At one later moment, its speed is 1.25 Mm/s (nonrelativistic because v is much less than the speed of light).

(a) Find the acceleration of the proton.

(b) Over what time interval does the proton reach this speed?

(c) How far does it move in this time interval?

(d) What is its kinetic energy at the end of this interval?
Answer #1

Solution :-

Given

Magnitude of electric field is , E = 615 N/C

The final speed is , v = 1.25 *10^6 m/s

Mass of proton , m= 1.67 *10^-27 kg

a)The force on the proton is F = qE

                                        ma = qE

Acceleration, a = (1.6*10^-19 C) * 615 N/C / (1.67 *10^-27 kg)

                         = 5.892*10^10 m/s^2

b) The time taken to reach this speed is

       t = v/a     [since v = at]

        t = (1.25 *10^6 m/s) /5.892 *10^10 m/s^2

        t = 2.1215 *10^-5 s

c) Distance moved in t is given by

   s = 1/2 at^2

   s = 1/2 * 5.892*10^10 m/s^2 *(2.1215*10^-5 s)^2

     = 13.26 m

d)The kinetic energy = 1/2 mv^2

                                 = 0.5 (1.67 *10^-27 kg)( 1.25 *10^6 m/s)^2

                                 = 1.305×10^-15 J